an object has a final velocity of 10m/s an initial velocity of 5m/s and an acceleration of 3m/s2.calculate the distance covered by this object in its motion.



1
Expert's answer
2022-01-02T18:24:18-0500
"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a},""d=\dfrac{(10\ \dfrac{m}{s})^2-(5\ \dfrac{m}{s})^2}{2\times3\ \dfrac{m}{s^2}}=12.5\ m."